Solve 3^x^+4 = 7^x^-1
Answer by lwsshak3(11628)

Solve 3^x^+4 = 7^x^-1
3^x^4 = 7^x^-1
take log of both sides
x^4*log3=x^-1*log7
x^4/x^-1=log7/log3
x^5=log7/log3
x=(log7/log3)^(1/5)
using calculator
x≈1.1211
Check: w/calculator
3^x^4=3^(1.1211^4)≈5.67...
7^x^-1=7^(1.1211^-1)≈5.67..
left side≈right side